# Language Arithmetics: Towards Systematic Language Neuron Identification and Manipulation

### v2

@@ Abstract
Through language arithmetics, i.e. systematic activation addition and multiplication, we steer models to deactivate unwanted languages and activate desired ones, outperforming simpler replacement approaches. These interventions effectively guide behavior across five multilingual tasks: language forcing, translation, QA, comprehension, and NLI. Manipulation is more successful for high-resource languages, while typological similarity improves effectiveness. We also demonstrate that cross-lingual neuron steering enhances downstream performance and reveal internal "fallback" mechanisms for language selection when neurons are progressively deactivated. Our code is made publicly available at https://github.com/d-gurgurov/Language-Neurons-Manipulation.
@@ 1 Introduction
The emergence of large language models (LLMs) with impressive multilingual capabilities has raised fundamental questions about how these systems internally represent and process different languages Wendler et al. (2024); Zhao et al. (2024). While models like Llama-3 Grattafiori et al. (2024) and Gemma-3 Team et al. (2024) perform well across dozens of languages despite limited multilingual training data, the neural mechanisms underlying this competence, such as the emergence of language-specific spaces, cross-lingual feature sharing, and internal language routing, are not fully understood. Understanding these mechanisms is crucial not only for advancing theoretical insights into multilingual representation learning, but also for building more controllable and interpretable language technologies Amodei et al. (2016); Gabriel (2020); Singh et al. (2024).
Figure 1: Success rates of language forcing when deactivating neurons for the input language and activating those of a target language for Llama-3.1-8B.
@@
In this work, we present the first large-scale investigation of language-specific neuron identification and manipulation in the Llama-3.1-8B Grattafiori et al. (2024), Mistral-Nemo-12B AI (2024), and Aya-Expanse-8B & 32B Dang et al. (2024) models across 21 typologically diverse languages. We examine the distribution of language-sensitive neurons, their cross-linguistic overlap, and the evolution of output language distributions across layers. Using the Language Activation Probability Entropy (LAPE) method (Tang et al., 2024), we identify neurons with strong language preferences and report several key findings. Unlike previously tested models such as Llama-2 Touvron et al. (2023), BLOOM Le Scao et al. (2023), and Phi-2 Javaheripi et al. (2023), the models we study exhibit concentrated language-specific activity in their deeper layers. Analyzing output language distributions with logit lens Wendler et al. (2024) across layers reveals that language generation predominantly occurs in the final layers, aligning with the concentration of language-sensitive neurons. Non-Latin script languages show both a higher number of specialized neurons and less cross-language overlap, while typologically related Germanic and Romance languages share more neurons–reflecting linguistic proximity and possibly orthographic similarities in their internal representations.
@@
Building on these findings, we establish simple additive and multiplicative neuron intervention techniques, which we call language arithmetics [^1] The naming follows conceptually similar works on task and prompt arithmetics Ilharco et al. (2023); Belanec et al. (2024)., to test the causal role of language-specific neurons in controlling language use while preserving internal representations. Unlike prior approaches that rely on activation replacement Tang et al. (2024); Kojima et al. (2024), our approach adds target-language patterns to hidden states, while simultaneously multiplying unwanted language patterns with 0. We demonstrate that language arithmetics outperforms both replacement-based methods Tang et al. (2024); Kojima et al. (2024) and the widely used DiffMean Marks and Tegmark (2023) approach on a novel language forcing task, where models are expected to answer questions in a specific target language without being explicitly prompted, with especially strong results for languages better represented in pretraining. We further evaluate additive language arithmetic on four downstream tasks–machine translation, question answering, natural language inference, and machine comprehension–and show that targeted neuron activation improves performance without any task-specific fine-tuning, contrasting with the findings of Mondal et al. (2025). For example, our intervention improves translation scores by up to 10% and enables precise cross-lingual transfer. Additionally, we find that when dominant language signals are suppressed, models fall back to the next most probable language, suggesting the presence of internal "fallback" mechanisms for language selection.
@@ 2 Related Work
Our work extends this line of research in two dimensions: (1) We analyze the internal structure of LLMs without relying on auxiliary models such as SAEs, and (2) we look at a wider range of more diverse languages. We conduct a comprehensive study of language-specific neurons in LLaMA-3.1-8B Grattafiori et al. (2024), Mistral-Nemo-12B AI (2024), and Aya-Expanse-8B & 32B Dang et al. (2024) across 21 languages using the LAPE method (Tang et al., 2024), systematically investigating both neuron identification and manipulation to assess their effectiveness for multilingual control and downstream performance improvement.
@@ 3.1 Identification Method
Language Activation Probability Entropy (LAPE) Tang et al. (2024) identifies language-specific neurons within LLMs by analyzing activation patterns across different languages in the FFN modules of a transformer-based language model. For each neuron, it measures how often that neuron activates when processing inputs in various languages. These activation frequencies are normalized into a probability distribution over languages, and entropy Shannon (2001) is used to quantify how language-specific each neuron’s behavior is.
Intuitively, neurons with low LAPE values are considered language-specific since their activation probabilities are concentrated on one or two languages, showing minimal activity for others. We select the bottom K% (K={1..5}) of neurons by LAPE score after filtering out neurons with weak overall activity: neurons are excluded if none of their language activations exceed the 95th percentile of all activations (filter rate). To ensure neurons exhibit clear language preferences, they further get refined by verifying that their activation probability for a given language surpasses the 95th percentile (activation threshold). These thresholds help eliminate noisy or non-informative neurons and retain only those with high, selective activation characteristics.
@@ 3.3 Models and Data
We use the Llama-3.1-8B base model Grattafiori et al. (2024) for all experiments. Notably, Llama-3.1 is a decoder-only model that was neither explicitly trained for multilingual tasks nor instruction-tuned. Its pre-training data includes approximately 5% non-English content spanning over 30 languages. We also include Mistral-Nemo Mistral AI Team (2024), a slightly larger 12-B-parameter base model, featuring strong multilingual, reasoning, and coding performance in its size class. Finally, we evaluate the Aya Expanse family Dang et al. (2024): Aya-Expanse-8B, an open-weight, instruction-tuned multilingual model optimized via data arbitrage, preference training, and model-merging; and its larger sibling, Aya-Expanse-32B–both supporting 23 languages and offering state-of-the-art multilingual performance.
For neuron identification, we use the CulturaX corpus Nguyen et al. (2023), a large-scale multilingual dataset comprising over 6.3 trillion tokens across 167 languages. The data is sourced from Common Crawl Wenzek et al. (2019) and Wikipedia Foundation and has undergone extensive cleaning and language identification to ensure high quality. To compute language-specific activations and obtain LAPE values, we truncate the dataset for each language to 500MB due to efficiency reasons. For per-layer output experiments, we use 6 questions that are available in all languages, as described in Section 4.2.
Figure 2: Neuron overlap between languages and language families in Llama-3.1, based on the top 1% of neurons identified as language-specific. Diagonals show counts per language; off-diagonals show overlaps. Asterisks mark non-Latin script languages.
@@ 3.4 Results
Language Neuron Distribution: Figure 3 and Appendix E show total neuron activation distributions with minimal early-layer activity and pronounced peaks in later layers across all 21 languages. Interestingly, language-specific processing in Llama-3.1, as well as the other tested models, demonstrates a strong concentration in layers 17-28, peaking at layers 27-28. Earlier Layers 0-15 show relatively minimal language-specific activity, with only modest peaks around layers 3-4 and 8-9. This contrasts with previous findings by Tang et al. (2024), Kojima et al. (2024), and Zhao et al. (2024), who report both early- and late-layer language neuron specialization.
Individual language distribution analysis (see Appendices E and F) reveals typologically coherent clustering, with Germanic languages (German (de), Dutch (nl), Afrikaans (af), Danish (da), Norwegian (no), Swedish (sv)) showing similar language neuron distribution patterns concentrated around layers 24-26. Tibetan (bo) and Chinese (zh) (the only Sino-Tibetan languages we tested) demonstrate the most concentrated deep-layer activation, while Romance languages display more varied patterns. These patterns are observed across all tested models.
Figure 3: Layer-wise distribution of language-specific neurons in Llama-3.1 across 21 languages. Similar patterns for individual languages and other models are shown in Appendix E.
Language Neuron Overlap: The language neuron overlap matrices in Figure 2 and Appendices E and F show that language families exhibit internal cohesion with substantial neuron overlap, as also observed by Tan et al. (2024). In LLama-3.1, for instance, Germanic languages show particularly high overlap values (e.g., Swedish (sv) and Danish (da): 290 neurons representing 76.3% and 74.7% of their respective language-specific neurons; Dutch (nl) and Afrikaans (af): 250 neurons representing 76.5% and 59.0% respectively), possibly confirming that the similar activation patterns observed in Figure 13 reflect shared neurons rather than coincidental distributions.
@@
Per-layer Language Analysis: For more English-centric models (see Appendix G) such as Llama-3.1 and Mistral-Nemo, we observe that the target language tokens predominantly emerge in the final layers (25–30), aligning with the concentration of language-specific neurons in those layers. English tokens slightly begin to appear in the initial layers and persist through to the output. This pattern partly aligns with findings by Wendler et al. (2024) and Schut et al. (2025). Language entropy in these models increases across layers: the model appears confident in its language prediction in early layers, but entropy rises from around layer 20 onward, suggesting that the decision about which language to generate is made in the final layers.
Figure 4: Predicted probabilities of target languages for Llama-3.1 using logit lens outputs from each layer and FastText for language identification. Results for other models are shown in Appendix G.
@@ 4.1 Method
Table 1: Overall success rates (%) of language forcing for Llama-3.1-8B using three intervention types and two manipulation strategies across different top-k% neuron thresholds.
@@ 4.2 Tasks and Data
First, we design a novel controlled language-forcing task using six simple questions ("How are you today?", "What is your name?", "What year is it now?", "What is your favorite color?", "What is the weather like?", "Where are you from?") translated into 21 languages using Google Translate Wu et al. (2016). For each input question, we deactivate the neurons associated with the source language by setting their activations to zero or a small negative value (see Appendix H) and activate the neurons associated with a desired target language. We then use FastText (Joulin et al., 2016) to identify the language of the generated output. This setup evaluates the model’s ability to override the input language and generate responses in the specified target language.
@@ 4.3 Results
Language Forcing: For the language forcing task, we compare two neuron manipulation strategies: (1) deactivating neurons associated with the source language and simultaneously activating neurons for the target language, and (2) only activating the neurons corresponding to the desired target language. Our results show that the first approach more reliably steers the model to generate output in the target language (Table 1 and Appendix H). We additionally compare our additive language arithmetic intervention with the simple replacement methods Tang et al. (2024); Kojima et al. (2024) and the widely used DiffMean approach Marks and Tegmark (2023); Panickssery et al. (2023) [^2] For DiffMean, we construct the steering vector as the difference between the mean activation for the target language (positive class) and that of all 20 other languages (negative class). (see Table 1 and Appendix H). We find that the additive method performs best, likely because it is less disruptive to the model’s internal representations.
@@
Figure 8: XNLI changes over the baseline (measured by accuracy score) when activating language-specific neurons for LLama-3.1 (4%).
@@
Overall, we observe modest improvements (up to 10–20%) or neutral effects for most language pairs when activating neurons from various languages, particularly within the same language family, in contrast to the negative transfer effects reported by Mondal et al. (2025). For instance, activating Romance or Germanic language neurons consistently benefits other languages from the same families across all tasks. Task type influences cross-linguistic effects: translation and question answering show more consistent positive transfer, while machine comprehension and natural language inference exhibit mixed results, including occasional negative transfer from distant languages. Activating the actual target language typically improves performance across all tasks and languages. Interestingly, some Romance and Germanic languages benefit even from activating Sino-Tibetan and Other language neurons–likely due to the larger number of neurons in those groups and their partial overlap with the other families at current activation ratios. Finally, Hindi and Urdu show disproportionate gains from nearly all language activations, possibly due to polysemantic neurons–units encoding multiple concepts (Elhage et al., 2022).
Language "Fallbacks": By progressively deactivating high-resource language neurons in Llama-3.1-8B (see Appendix I), we observe that the model exhibits internal "fallback" strategies for language generation. When English (en) neurons are deactivated, the model defaults to Spanish (es), French (fr), and Russian (ru) in a few cases. Deactivating English and French shifts responses primarily to Russian (ru), with some in Italian (it), Spanish (es), and Arabic (ar). Further removal of Latin-script languages–Spanish, Italian, and Portuguese (pt)–promotes generation in Russian, Arabic, and Chinese. Deactivating German, Russian, and Arabic leads to "fallback" responses in Thai, Vietnamese, and Chinese. Notably, some English, Spanish, and French responses persist despite deactivation, reflecting the model’s strong priors for these languages. Similar patterns are observed for Mistral-Nemo, but with a slightly different language hierarchy. We leave a more detailed investigation of these dynamics to future work.

### v3

@@ Abstract
Through language arithmetics, i.e. systematic activation addition and multiplication, we steer models to deactivate unwanted languages and activate desired ones, outperforming established replacement approaches. These interventions effectively guide behavior across five multilingual tasks: language forcing, translation, QA, comprehension, and NLI. Manipulation is more successful for high-resource languages, while typological similarity improves effectiveness. We also demonstrate that neuron steering enhances downstream performance and reveal internal "fallback" mechanisms for language selection when neurons are progressively deactivated. Our code is made publicly available at https://github.com/d-gurgurov/Language-Neurons-Manipulation.
@@ 1 Introduction
The emergence of large language models (LLMs) with impressive multilingual capabilities has raised fundamental questions about how these systems internally represent and process different languages Wendler et al. (2024); Zhao et al. (2024). While models like Llama-3 Grattafiori et al. (2024) and Gemma-3 Team et al. (2024) perform well across dozens of languages despite limited multilingual training data, the neural mechanisms underlying this competence are not fully understood. Understanding these mechanisms is crucial not only for advancing theoretical insights into multilingual representation learning, but also for building more controllable and interpretable language technologies Amodei et al. (2016); Gabriel (2020); Singh et al. (2024).
Figure 1: Success rates of language forcing when deactivating neurons for the input language and activating those of a target language for Llama-3.1-8B. The input question is presented in the language corresponding to the deactivated neurons. Top 5% of neurons are considered.
@@
In this work, we present the first large-scale investigation of language-specific neuron identification and manipulation in the Llama-3.1-8B Grattafiori et al. (2024), Mistral-Nemo-12B AI (2024), and Aya-Expanse-8B & 32B Dang et al. (2024) models across 21 typologically diverse languages. We examine the distribution of language-sensitive neurons, their cross-linguistic overlap, and the evolution of output language distributions across layers. Using the Language Activation Probability Entropy (LAPE) method (Tang et al., 2024), we identify neurons with strong language preferences and report several key findings. Unlike previously tested models such as Llama-2 Touvron et al. (2023), BLOOM Le Scao et al. (2023), and Phi-2 Javaheripi et al. (2023), the models we study exhibit concentrated language-specific activity in their deeper layers. Analyzing output language distributions with logit lens Nostalgebraist (2020) across layers reveals that language generation predominantly occurs in the final layers, aligning with the concentration of language-sensitive neurons. Non-Latin script languages show both a higher number of specialized neurons and less cross-language overlap, while typologically related Germanic and Romance languages share more neurons–reflecting linguistic proximity and possibly orthographic similarities in their internal representations.
@@
Building on these findings, we establish simple additive and multiplicative neuron intervention techniques, which we call language arithmetics [^1] The naming follows conceptually similar works on task and prompt arithmetics Ilharco et al. (2023); Belanec et al. (2024)., to test the causal role of language-specific neurons in controlling language use while preserving internal representations. Unlike prior approaches that rely on activation replacement Tang et al. (2024); Kojima et al. (2024), our approach adds target-language patterns to hidden states, while simultaneously multiplying unwanted language patterns with 0. We demonstrate that language arithmetics outperforms both replacement-based methods Tang et al. (2024); Kojima et al. (2024) and the widely used DiffMean Marks and Tegmark (2023) approach on a novel language forcing task, where models are expected to answer questions in a specific target language without being explicitly prompted, with especially strong results for languages better represented in pre-training. We further evaluate additive language arithmetic on four downstream tasks–machine translation, question answering, natural language inference, and machine comprehension–and show that targeted neuron activation improves performance without any task-specific fine-tuning, contrasting with the findings of Mondal et al. (2025). For example, our intervention improves translation scores by up to 10% and enables cross-lingual transfer. Additionally, we find that when dominant language signals are suppressed, models fall back to the next most probable language, suggesting the presence of internal "fallback" mechanisms for language selection.
@@ 2 Related Work
Our work extends this line of research along two key dimensions: (1) we analyze the internal structure of LLMs directly, without relying on auxiliary models such as SAEs, and (2) we examine a broader and more diverse set of languages. Specifically, we conduct a comprehensive study of language-specific neurons in Llama-3.1-8B Grattafiori et al. (2024), Mistral-Nemo-12B AI (2024), and Aya-Expanse-8B & 32B Dang et al. (2024) across 21 languages. Our analysis employs the LAPE method (Tang et al., 2024), grounded in information theory, and systematically investigates both neuron identification and manipulation to evaluate their effectiveness for multilingual control and downstream performance improvement.
@@ 3.1 Identification Method
Language Activation Probability Entropy (LAPE) Tang et al. (2024) identifies language-specific neurons within LLMs by analyzing activation patterns across different languages in the FFN modules of a transformer-based language model.
For each neuron $j$ in layer $i$, the activation probability when processing text in language $k$ is computed as:
$p^{k}_{i,j}=\mathbb{E}[\mathbf{I}(\text{act\_fn}(\tilde{h}_{i}W^{i}_{1})_{j}>0)\mid\text{language }k]$
where $\mathbf{I}$ is the indicator function. The activation probabilities across all languages form a distribution $p_{i,j}=(p^{1}_{i,j},p^{2}_{i,j},\ldots,p^{l}_{i,j})$, which is L1-normalized to obtain $p^{\prime}_{i,j}$. LAPE is then calculated using Shannon entropy Shannon (2001):
$\text{LAPE}_{i,j}=-\sum_{k=1}^{l}p^{\prime k}_{i,j}\log(p^{\prime k}_{i,j})$
Intuitively, neurons with low LAPE values are considered language-specific since their activation probabilities are concentrated on one or two languages, showing minimal activity for others.We select the bottom K% (K={1..5}) of neurons by LAPE score as candidate language-specific neurons through a three-step filtering process: (1) we exclude neurons with weak overall activity where none of their language activations exceed the 95th percentile (filter rate = 0.95), (2) from the remaining neurons, we select the bottom K% by LAPE score to identify those with the most language-specific activation patterns, and (3) we assign each selected neuron to all languages where its activation probability $p^{k}_{i,j}$ exceeds the 95th percentile threshold (activation threshold = 0.95). While neurons can be assigned to multiple languages if they show strong activation across them, low LAPE scores typically result in assignments to only one or two languages.
@@ 3.3 Models and Data
We use the Llama-3.1-8B base model Grattafiori et al. (2024) for all experiments. Notably, Llama-3.1 is a decoder-only model that was neither explicitly trained for multilingual tasks nor instruction-tuned. Its pre-training data includes approximately 5% non-English content spanning over 30 languages. We also include Mistral-Nemo Mistral AI Team (2024), a slightly larger 12-B‑parameter base model, featuring strong multilingual, reasoning, and coding performance in its size class. Finally, we evaluate the Aya Expanse family Dang et al. (2024): Aya-Expanse-8B, an open‑weight, instruction‑tuned multilingual model optimized via data arbitrage, preference training, and model‑merging; and its larger sibling, Aya-Expanse-32B–both supporting 23 languages and offering state‑of‑the‑art multilingual performance.
For neuron identification, we use the CulturaX corpus Nguyen et al. (2023), a large-scale multilingual dataset comprising over 6.3 trillion tokens across 167 languages. The data is sourced from Common Crawl Wenzek et al. (2019) and Wikipedia Foundation and has undergone extensive cleaning and language identification to ensure high quality. To compute language-specific activations and obtain LAPE values, we truncate the dataset for each language to 500MB (approx. 100M tokens) due to efficiency reasons. Our experiments focus on 21 representative languages, detailed in Appendix B. For per-layer output experiments, we use 6 questions that are available in all languages, as described in Section 4.2.
Figure 2: Neuron overlap between languages and language families in Llama-3.1-8B, based on the top 1% of neurons identified as language-specific. Diagonals show counts per language; off-diagonals show overlaps. Asterisks mark non-Latin script languages.
@@ 3.4 Results
Language Neuron Distribution: Figure 3 and Appendix D show total neuron activation distributions with minimal early-layer activity and pronounced peaks in later layers across all 21 languages. Interestingly, language-specific processing in Llama-3.1, as well as the other tested models, demonstrates a strong concentration in layers 17-28, peaking at layers 27-28. Earlier Layers 0-15 show relatively minimal language-specific activity, with only modest peaks around layers 3-4 and 8-9. This contrasts with previous findings by Tang et al. (2024), Kojima et al. (2024), and Zhao et al. (2024), who report both early- and late-layer language neuron specialization.
Individual language distribution analysis (see Appendix D) reveals typologically coherent clustering, with Germanic languages (German (de), Dutch (nl), Afrikaans (af), Danish (da), Norwegian (no), Swedish (sv)) showing similar language neuron distribution patterns concentrated around layers 24-26. Tibetan (bo) and Chinese (zh) (the only Sino-Tibetan languages we tested) demonstrate the most concentrated deep-layer activation, while Romance languages display more varied patterns. These patterns are observed across all tested models.
Figure 3: Layer-wise distribution of language-specific neurons for individual languages in Llama-3.1-8B. Other models exhibit similar patterns (Appendix D).
Language Neuron Overlap: The language neuron overlap matrices in Figure 2 and Appendix D show that language families exhibit internal cohesion with substantial neuron overlap, as also observed by Tan et al. (2024). In Llama-3.1, for instance, Germanic languages show particularly high overlap values (e.g., Swedish (sv) and Danish (da): 290 neurons representing 76.3% and 74.7% of their respective language-specific neurons; Dutch (nl) and Afrikaans (af): 250 neurons representing 76.5% and 59.0% respectively), possibly confirming that the similar activation patterns observed in Figure 19 of Appendix D reflect shared neurons rather than coincidental distributions.
@@
Per-layer Language Analysis: For more English-centric models (see Appendix E) such as Llama-3.1 and Mistral-Nemo, we observe that the target language tokens predominantly emerge in the final layers (25–30), aligning with the concentration of language-specific neurons in those layers. English tokens slightly begin to appear in the initial layers and persist through to the output. This pattern partly aligns with findings by Wendler et al. (2024) and Schut et al. (2025). Language entropy in these models increases across layers: the model appears confident in its language prediction in early layers, but entropy rises from around layer 20 onward, suggesting that the decision about which language to generate is made in the final layers.
Figure 4: Predicted probabilities of target languages for Llama-3.1 using logit lens outputs from each layer and FastText for language identification. Results for other models are shown in Appendix E.
@@ 4.1 Method
Table 1: Overall success rates (%) of language forcing for Llama-3.1-8B using three intervention types and two manipulation strategies across different top-k% neuron thresholds. The results for Mistral-Nemo and Aya-Expanse follow similar patterns (see Appendix F).
@@ 4.2 Tasks and Data
First, we design a novel controlled language-forcing task using six simple questions ("How are you today?", "What is your name?", "What year is it now?", "What is your favorite color?", "What is the weather like?", "Where are you from?") translated into 20 languages using Google Translate Wu et al. (2016). For each input question, we deactivate the neurons associated with the source language by setting their activations to zero (see Appendix F) and activate the neurons associated with a desired target language by applying the additive intervention. We then use FastText (Joulin et al., 2016) to identify the language of the generated output. This setup evaluates the model’s ability to override the input language and generate responses in the specified target language.
@@ 4.3 Results
Language Forcing: For the language forcing task, we compare two neuron manipulation strategies: (1) deactivating neurons associated with the source language and simultaneously activating neurons for the target language, and (2) only activating the neurons corresponding to the desired target language. Our results show that the first approach more reliably steers the model to generate output in the target language (Table 1 and Appendix F). We additionally compare our additive language arithmetic intervention with the simple replacement methods Tang et al. (2024); Kojima et al. (2024) and the widely used DiffMean approach Marks and Tegmark (2023); Panickssery et al. (2023) [^2] For DiffMean, we construct the steering vector as the difference between the mean activation for the target language (positive class) and that of all 20 other languages (negative class). (see Table 1 and Appendix F). We find that the additive method performs best across all tested models, likely because it is less disruptive to the model’s internal representations.
@@
Figure 8: XNLI changes over the baseline (measured by accuracy score) when activating language-specific neurons for Llama-3.1 (5%).
@@
Figure 9: Progressive deactivation of language-specific neurons for high-resource languages in Llama-3.1-8B. In this setting, language neurons are deactivated by setting their activations to 1-1; using 0 does not yield the desired effect. After each deactivation round, the model is prompted with 70 English questions, and we then identify the output language of its responses.
Activating target language neurons improves performance in nearly all cases across all tasks and languages, with gains typically ranging from 5-15%, as opposed to the negative effects reported by Mondal et al. (2025). Cross-lingual transfer effects vary substantially by task type. For generative tasks–FLORES translation and XQuAD question answering–we observe minimal cross-lingual improvements and frequent performance degradations when activating non-target languages. This pattern is particularly pronounced in FLORES, where activating neurons from different languages often causes the model to generate outputs in the activated language rather than the target language, explaining the consistent negative transfer. XQuAD exhibits similar behavior, though somewhat less severe, as the generative nature of the task makes it susceptible to language confusion. In contrast, discriminative tasks–Belebele reading comprehension and XNLI natural language inference–show markedly different patterns. Here, activating target language neurons remains beneficial, but we also observe substantial positive cross-lingual transfer, particularly from related languages within the same family. Romance language activations benefit other Romance languages, and Germanic activations similarly help Germanic languages, with improvements of 5-10% in many cases. Notably, Hindi and Urdu demonstrate exceptional robustness, maintaining or improving performance under nearly all language activations across both task types, possibly due to polysemantic neurons, units encoding multiple concepts (Elhage et al., 2022).
Language "Fallbacks": By progressively deactivating high-resource language neurons in Llama‑3.1‑8B (see Figure 9 and Appendix G), we observe that the model exhibits internal "fallback" strategies for language generation. When English (en) neurons are deactivated, the model defaults to Spanish (es), French (fr), and Russian (ru) in a few cases. Deactivating English and French shifts responses primarily to Russian (ru), with some in Italian (it), Spanish (es), and Arabic (ar). Further removal of Latin-script languages–Spanish, Italian, and Portuguese (pt)–promotes generation in Russian, Arabic, and Chinese. Deactivating German, Russian, and Arabic leads to "fallback" responses in Thai, Vietnamese, and Chinese. Notably, some English, Spanish, and French responses persist despite deactivation, reflecting the model’s strong priors for these languages. Similar patterns are observed for Mistral-Nemo, but with a slightly different language hierarchy. We leave a more detailed investigation of these dynamics to future work.